Isabella is making a quilt She takes one square piece of cloth, folds it in half and cuts along the fold making to rectangles. The perimeter of each retangle is 27 cms. What is the perimeter of the square piece of cloth? How do I figure this out by not know the dimensions of the rectangle.
Answer by checkley79(3341)

The length of the half square is x & the width is x/2.
2l+2w=perimeter.
2x+2(x/2)=27
2x+x=27
3x=27
x=27/3
x=9 cm. is the length of the original square.
4*9=36 cm. is the perimeter of the original square.